Find the roots of the following quadratic equations, if they exist, by the method of completing the square: 
 4x^2 + 4√3x + 3 = 0

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
Converting the equation into a^2+2ab+b^2form
⇒ (2x)^2 + 2 × 2x × √3 + (√3) 2 = 0
⇒ (2x + √3) 2 = 0
⇒ (2x + √3) = 0 and (2x + √3) = 0
either x = -√3/2 or x = -√3/2.

Thanks